# LLMs generate structurally realistic social networks but overestimate political homophily

### v1

@@ Abstract
Generating social networks is essential for many applications, such as epidemic modeling and social simulations. Prior approaches either involve deep learning models, which require many observed networks for training, or stylized models, which are limited in their realism and flexibility. In contrast, LLMs offer the potential for zero-shot and flexible network generation. However, two key questions are: (1) are LLM’s generated networks realistic, and (2) what are risks of bias, given the importance of demographics in forming social ties? To answer these questions, we develop three prompting methods for network generation and compare the generated networks to real social networks. We find that more realistic networks are generated with “local” methods, where the LLM constructs relations for one persona at a time, compared to “global” methods that construct the entire network at once. We also find that the generated networks match real networks on many characteristics, including density, clustering, community structure, and degree. However, we find that LLMs emphasize political homophily over all other types of homophily and overestimate political homophily relative to real-world measures.
LLMs generate structurally realistic social networks
but overestimate political homophily
Serina Chang1,*, Alicja Chaszczewicz1,*, Emma Wang1, Maya Josifovska1,2, Emma Pierson3, Jure Leskovec1 1Department of Computer Science, Stanford University 2Department of Computer Science, University of California, Los Angeles 3Department of Computer Science, Cornell University *These authors contributed equally.
@@ 1 Introduction
The ability to generate realistic social networks is crucial for many applications, when the true social network cannot be observed (e.g., for privacy reasons) or a realistic network is desired between hypothetical individuals. For example, in epidemic modeling, synthetic social networks are frequently used so that researchers can model the spread of disease based on who has come into contact with whom Barrett et al. (2009); Block et al. (2020). Synthetic networks are also useful for simulating and analyzing social media platforms Pérez-Rosés and Sebé (2015); Sagduyu et al. (2018) and social phenomena, such as polarization and opinion dynamics Dandekar et al. (2013); Das et al. (2014).
@@
Deep learning approaches to network generation typically require training on many domain-specific networks You et al. (2018), making it difficult to generalize to new settings where networks are not yet observed. Classical models for network generation require far less training, but these stylized models make rigid and unrealistic assumptions about how networks form. For example, Erdős–Rényi models assume that each edge forms with a uniform probability $p$ Erdös and Rényi (1959). More realistic models, like small-world models Watts and Strogatz (1998) or stochastic block models Holland et al. (1983), are still limited by a predefined, small set of numerical hyperparameters, missing the full complexity of real social interactions.
@@
In contrast, generating social networks with large language models (LLMs) has the potential to address these limitations. LLMs possess zero-shot capabilities, enabling network generation without training. LLMs can also generate networks in a flexible manner, based on natural language descriptions of each person in the network. A key question, however, is whether LLMs can generate realistic social networks. On one hand, LLMs have demonstrated capabilities to realistically simulate human responses and interactions Aher et al. (2023); Park et al. (2023); Argyle et al. (2023), suggesting that they may be able to generate realistic social networks as well. On the other hand, LLMs sometimes struggle with reasoning over graphs Wang et al. (2023); Fatemi et al. (2024) and it is unclear if their language abilities generalize to structured objects like networks, so that they can reproduce structural characteristics of social networks such as low density and long-tailed degree distributions.
@@
Furthermore, a central concern with using LLMs in social settings is bias. Prior works have shown that LLMs produce stereotyped descriptions of individuals based on their demographics Cheng et al. (2023a, b) and skew towards the liberal opinions Santurkar et al. (2023). These demographics, such as gender and political affiliation, play essential roles in the formation of real-world social networks, resulting in well-documented demographic homophily McPherson et al. (2001); Kossinets and Watts (2009). Thus, we cannot evaluate whether LLMs’ social network generation is realistic without incorporating demographics into our experiments; at the same time, we need to analyze how LLMs reason about these demographic features and investigate potential signs of bias.
@@
To answer these questions, we propose three prompting methods for social network generation (Figure 1). First, we find that “local” methods, where the LLM takes on the perspective of one person at a time, yield more realistic networks than “global” methods, where the LLM constructs the entire network at once, even though the LLM receives less information in the local setting. The LLM is able to generate networks that match real networks on many structural characteristics, including density, clustering, connectivity, and degree distribution. The LLM also exhibits clear demographic homophily, across gender, age, race/ethnicity, religion, and political affiliation. However, the LLM consistently emphasizes homophily in political affiliation above all other demographic variables, and overestimates levels of political homophily compared to reported levels in real online and offline social networks. Finally, we find that incorporating LLM-generated interests does not reduce political homophily, since the interests themselves encode political stereotypes. Overall, our work demonstrates the promise of using LLMs for social network generation while calling attention to challenges around integrating demographics. [^1] Our code is available at https://github.com/snap-stanford/llm-social-network, along with our generated personas and generated social networks.
@@ 2 Related work
Prior work has demonstrated LLMs’ abilities to realistically simulate human responses and interactions Aher et al. (2023); Park et al. (2023); Argyle et al. (2023). However, while simulating interactions over networks, existing work focuses less on using LLMs to generate the networks themselves, either making simplistic assumptions about the network structure such as sampling agents randomly to interact Park et al. (2022); Chuang et al. (2023) or requiring human involvement in building the network Gao et al. (2023); Zhou et al. (2024). To improve the realism and usability of these simulations, it is essential to also explore LLMs’ abilities to generate the network structure, a prerequisite to simulating interactions over networks.
A few contemporaneous works have explored LLMs for social network generation, with different focuses from ours. Marzo et al. (2023) focus on degree distribution, showing that scale-free networks emerge from interactions between LLMs. He et al. (2023) focus on content homophily, analyzing a simulated society powered by LLM chatbots. The most similar work to ours is Papachristou and Yuan (2024), who analyze whether LLMs demonstrate network formation principles, such as preferential attachment and homophily. While their work establishes the existence of general network principles, our work compares generated networks to real social networks directly, computing many network metrics, and shows that all metrics can be matched at once, while their experiments primarily explore one principle at a time, with a different prompt for each principle. Also, to test homophily, they consider hobby, favorite color, and location, while we explore key demographic features.
@@
Using LLMs in social contexts raises concerns of biases and stereotyping Cheng et al. (2023a, b); Wang et al. (2024). When responding to public opinion or political questions, LLMs’ answers skew liberal Santurkar et al. (2023); Hartmann et al. (2023). When assigned a persona, LLMs show worse reasoning capabilities when assigned certain demographics Gupta et al. (2024) and produce more toxic content under certain personas Deshpande et al. (2023). However, bias in the context of social network generation remains unexplored. In this work, we investigate such biases by studying the effects of demographic variables on LLM-generated social networks.
Figure 1: Our three prompting methods to generate social networks with LLMs. See full prompts in Figures 17-19.
@@
Classical models for network generation do not require much training, but their simplicity often results in unrealistic network structures. For example, Erdős–Rényi models assume that each edge forms with uniform probability Erdös and Rényi (1959), Watts–Strogatz models Watts and Strogatz (1998) generate small-world networks but struggle to produce realistic degree distributions, and Barabási–Albert models Barabási and Albert (1999) capture power-law degree distributions but miss community structure and clustering. In contrast, machine learning approaches aim to learn graph generation directly from observed data, resulting in more realistic generated networks You et al. (2018); Simonovsky and Komodakis (2018); Guo and Zhao (2023). However, these models require a set of training networks, and typically have fixed feature sets, so they cannot be applied in situations without access to real-world network data, and cannot be easily extended to incorporate new features, while LLMs can be.
@@
The use of LLMs has been explored for various graph tasks Li et al. (2024); Jin et al. (2024), such as graph reasoning Wang et al. (2023); Fatemi et al. (2024), node classification Zhao et al. (2023); Chen et al. (2024); Ye et al. (2024), or tasks on knowledge graphs Pan et al. (2024). These works have shown that LLMs possess preliminary graph reasoning capabilities, but struggle on larger graphs or harder tasks, such as finding a Hamilton path Wang et al. (2023). Yao et al. (2024) is one of the first to explore LLMs for graph generation, but they explore it in the context of molecule generation, not social networks.
@@ 3 Generating social networks
Our process for generating social networks involves two steps: first, constructing a set of personas, and second, having the LLM generate social networks over those personas. In this section, we summarize the two steps, with details in Appendix C.
@@ 3.1 Persona construction
For each persona, we include their gender, age, race/ethnicity, religion, and political affiliation, which are salient dimensions of homophily in real social networks McPherson et al. (2001); Halberstam and Knight (2016). We sample these characteristics based on the distribution of the US population. Using US Census data US Census Bureau (2023), we acquire the joint distribution for gender, age, and race/ethnicity. Then, we sample the persona’s religion, conditioned on their race/ethnicity Statista (2016); PRRI Staff (2021), and political affiliation, conditioned on their gender and race/ethnicity Pew Research Center (2024); Sanchez and Foxworth (2022). In Section 5.3, we also experiment with including interests for each persona, instead of only demographic variables.
@@ 3.2 Network generation
In our first method, which we call “Global”, we provide the LLM with the entire list of personas, and prompt it to construct the network between them, in the form of edge pairs.
@@
In our second method, which we call “Local”, we have the LLM take on the identity of one persona at a time, e.g., by saying, “You are a Woman, Asian, age 54, Hindu, Independent.” We provide the LLM with the list of all other personas (in the same format) and prompt it to pick friends for the persona it is currently assigned. To construct the entire network, we iterate through all personas in a random order, and we keep an edge between personas A and B if the LLM selects B when acting as A or vice versa (so we do not require both to select each other). This method is inspired by techniques in machine learning that similarly model the graph generation process by iterating through nodes and selecting edges for each node at a time You et al. (2018).
@@
In our third method, we also assign the LLM one persona at a time, but in addition to providing the list of all other personas, we also provide information about the constructed graph so far. We experiment with providing each persona’s full list of current friends vs only their degree (i.e., total number of friends). These variations are similar to the preferential attachment experiments in Papachristou and Yuan (2024), where they also experimented with providing neighborhood information vs. only degree, although their experiments only considered the network and no demographic features. In contrast, our experiments—and in particular, the comparison of the Local to Sequential methods—reveal how providing demographic and network information compare to only providing demographic information.
@@ 4 Comparison to real networks
To evaluate the realism of our generated networks, we gather a set of real social networks from the CASOS [^2] http://www.casos.cs.cmu.edu/tools/datasets/external/index.php and KONECT [^3] http://konect.uni-koblenz.de/networks/ repositories. We kept networks that described social relations between individuals, which filtered out other types of networks, such as work-related interactions or visiting ties between families. We included eight real networks, which capture social ties within diverse communities, including physicians, students, and prisoners (see Appendix C.3 for details).
We extract graph-level and node-level metrics from the real networks and our generated networks, and compare their distributions. For consistency, we treat all networks as undirected. Since the number of nodes varies across networks, we focus on network metrics that are comparable across graphs of different sizes, and scale those that are dependent on network size based on how they are expected to scale in an Erdös-Rényi random graph Erdös and Rényi (1959). Below, we define and motivate the network metrics that we evaluate on.
@@
A basic property of a network is its density of edges, and social networks tend to be sparse, meaning lower density Wong et al. (2006). Density computed as the number of observed edges divided by the total number of possible edges in the network, which comes out to $\frac{2E}{N(N-1)}$, where $N$ is the number of nodes and $E$ is the number of edges in the network.
@@
Social networks are known to exhibit clustering, where one’s friends are likely to be friends with each other Alizadeh (2017). For a node $i$, its clustering coefficient is $\frac{2E_{i}}{N_{i}(N_{i}-1)}$, where $N_{i}$ is its number of neighbors and $E_{i}$ is the number of edges between its neighbors. The average clustering coefficient computes the average over nodes.
Proportion of nodes in the largest connected component (LCC).
Social networks are known to be well-connected Ugander et al. (2011), with the vast majority (over 99%) of the nodes in the largest connected component (LCC), i.e., the largest subgraph where all nodes within the subgraph are reachable by each other. Thus, as a metric, we compute the proportion of all nodes in the network that are in the LCC, $\frac{N_{LCC}}{N}$.
Average shortest path in LCC.
Social networks are not only well-connected, meaning nodes can reach each other, but also they can reach each other in relatively short paths Watts and Strogatz (1998); Alizadeh (2017). So, we measure the average shortest path over all pairs of nodes in the LCC, divided by $\log N_{LCC}$, since shortest paths scale with $\log N$ in Erdös-Rényi graphs Watts and Strogatz (1998). We compute shortest paths within the LCC instead of the entire network, since the distance between two disconnected nodes is infinite.
@@
Social networks exhibit strong community structure, with more edges within communities and fewer edges across communities Newman (2004). To measure community structure, first we use the Louvain algorithm Blondel et al. (2008) to partition the network into communities, then we assess the quality of the partition with modularity (Eq. 3). Higher levels of community structure correspond to higher modularity.
@@
Social networks are said to follow a power law degree distribution, where $P(k)\sim k^{-\gamma}$, for degree $k$ and constant $\gamma$ Barabási and Albert (1999). This results in long-tailed degree distributions with a few people having far more friends than most others. To measure degree distribution, we compute the degree of each node in the network, and to make degree comparable across graphs, we divide degree by $N$. To summarize degree distribution from a set of networks (e.g., all real networks), we pool all of the (normalized) degrees of the nodes in the networks in the set, and compute the distribution over the pooled degrees in bins of 0.05 (from 0 to 1).
@@
Finally, social networks are known to exhibit homophily, where “birds of a feather flock together”, i.e., people with similar traits are likelier to be friends McPherson et al. (2001); Kossinets and Watts (2009). To measure homophily, we consider the observed proportion of same-group edges (e.g., same gender), divided by the expected proportion of same-group edges, given the number of nodes in each group:
$H$ $=\frac{S_{\textrm{obs}}}{S_{\textrm{exp}}}=\frac{\frac{\sum_{ij}A_{ij}\cdot\mathbbm{1}[g_{i}=g_{j}]}{E}}{\frac{\sum_{g}N_{g}(N_{g}-1)}{N(N-1)}},$ (1)
where $A_{ij}$, as the adjacency matrix, is 1 if nodes $i$ and $j$ are connected and 0 otherwise; $g_{i}$ indicates node $i$’s group; and $N_{g}$ is the number of nodes in group $g$. Higher ratios (above 1) indicate homophily and lower ratios (below 1) indicate heterophily, meaning there are fewer same-group relations than expected (this appears, for example, in heterosexual dating networks).

### v2

@@ Abstract
Generating social networks is essential for many applications, such as epidemic modeling and social simulations. The emergence of generative AI, especially large language models (LLMs), offers new possibilities for social network generation: LLMs can generate networks without additional training or need to define network parameters, and users can flexibly define individuals in the network using natural language. However, this potential raises two critical questions: 1) are the social networks generated by LLMs realistic, and 2) what are risks of bias, given the importance of demographics in forming social ties? To answer these questions, we develop three prompting methods for network generation and compare the generated networks to a suite of real social networks. We find that more realistic networks are generated with “local” methods, where the LLM constructs relations for one persona at a time, compared to “global” methods that construct the entire network at once. We also find that the generated networks match real networks on many characteristics, including density, clustering, connectivity, and degree distribution. However, we find that LLMs emphasize political homophily over all other types of homophily and significantly overestimate political homophily compared to real social networks.
Codehttps://github.com/snap-stanford/llm-social-network
@@ 1 Introduction
The ability to generate realistic social networks is crucial for many applications, when the true social network cannot be observed (e.g., for privacy reasons) or a realistic network is desired between hypothetical individuals. For example, in epidemic modeling, synthetic social networks are frequently used so that researchers can model the spread of disease based on who has come into contact with whom (Barrett et al. 2009; Block et al. 2020). Synthetic networks are also useful for simulating and analyzing social media platforms (Pérez-Rosés and Sebé 2015; Sagduyu, Grushin, and Shi 2018) and social phenomena, such as polarization and opinion dynamics (Dandekar, Goel, and Lee 2013; Das, Gollapudi, and Munagala 2014).
@@
Deep learning approaches to network generation typically require training on many domain-specific networks (You et al. 2018), making it difficult to generalize to new settings where networks are not yet observed. On the other hand, some classical network models require far less or no training, since they only have a few parameters, but in exchange for this simplicity, they make rigid and unrealistic assumptions about how social networks form. For example, Erdős–Rényi models only take in two inputs, $n$ (the number of nodes) and $p$, and they assume each edge forms with uniform probability $p$ (Erdős and Rényi 1959).
@@
In contrast, LLMs balance these challenges: they can generate social networks in a zero-shot fashion, without any additional training or need to define network parameters, but they can also take in rich, flexible inputs describing individuals in natural language and use those inputs to generate a network between those individuals. A key question, however, is whether the social networks generated by LLMs are realistic. On one hand, LLMs have demonstrated capabilities to realistically simulate human responses and interactions (Aher, Arriaga, and Kalai 2023; Park et al. 2023; Argyle et al. 2023), suggesting that they may be able to generate realistic social networks as well. On the other hand, LLMs sometimes struggle with reasoning over graphs (Wang et al. 2023; Fatemi, Halcrow, and Perozzi 2024) and it is unclear if their language abilities generalize to structured objects like networks, so that they can reproduce structural characteristics of social networks such as low density and long-tailed degree distributions.
@@
Furthermore, a central concern with using LLMs in social settings is bias. Prior works have shown that LLMs produce stereotyped descriptions of individuals based on their demographics (Cheng, Durmus, and Jurafsky 2023; Cheng, Piccardi, and Yang 2023) and skew towards the liberal opinions (Santurkar et al. 2023). These demographics, such as gender and political affiliation, play essential roles in the formation of real-world social networks, resulting in well-documented demographic homophily (McPherson, Smith-Lovin, and Cook 2001; Kossinets and Watts 2009; Halberstam and Knight 2016). Thus, we cannot evaluate whether LLMs’ social network generation is realistic without incorporating demographics into our experiments; at the same time, we need to analyze how LLMs reason about these demographic features and investigate potential signs of bias.
@@
To answer these questions, we propose three zero-shot prompting methods for social network generation (Figure 1). First, we find that “local” methods, where the LLM takes on the perspective of one person at a time, yield more realistic networks than “global” methods, where the LLM constructs the entire network at once, even though the LLM receives less information in the local setting. The LLM is able to generate networks that match real networks on many structural characteristics, including density, clustering, connectivity, and degree distribution. The LLM also exhibits clear demographic homophily, across gender, age, race/ethnicity, religion, and political affiliation. However, the LLM consistently emphasizes homophily in political affiliation above all other demographic variables, and overestimates levels of political homophily compared to reported levels in real online and offline social networks. Finally, we find that incorporating LLM-generated interests does not reduce political homophily, since the interests themselves encode political stereotypes. Overall, our work demonstrates the promise of using LLMs for social network generation while calling attention to challenges around integrating demographics.
@@ 2 Related Work
Prior work has demonstrated LLMs’ abilities to realistically simulate human responses and interactions (Aher, Arriaga, and Kalai 2023; Park et al. 2023; Argyle et al. 2023) and studied the dynamics of LLM agents interacting in a population, such as how conventions or consensus arise (Ashery, Aiello, and Baronchelli 2024; Marzo, Castellano, and Garcia 2024). However, while simulating interactions over social networks, existing work focuses less on using LLMs to generate the networks themselves, either making simplistic assumptions about the network structure such as sampling agents randomly to interact (Park et al. 2022; Chuang et al. 2023; Ashery, Aiello, and Baronchelli 2024) or assuming fully connected networks (Marzo, Castellano, and Garcia 2024), or requiring human involvement in building the network (Gao et al. 2023; Zhou et al. 2024). To improve the realism and usability of these simulations, it is essential to also explore LLMs’ abilities to generate the network structure, a prerequisite to simulating interactions over networks.
A few contemporaneous works have explored LLMs for social network generation, with different focuses from ours. Marzo, Pietronero, and Garcia (2023) focus on degree distribution, showing that scale-free networks emerge from interactions between LLMs. He et al. (2023) focus on content homophily, analyzing a simulated society powered by LLM chatbots. The most similar work to ours is Papachristou and Yuan (2024), who analyze whether LLMs demonstrate network formation principles, such as preferential attachment and homophily. While their work establishes the existence of general network principles, our work compares generated networks to real social networks directly, computing many network metrics, and shows that all metrics can be matched at once, while their experiments primarily explore one principle at a time, with a different prompt for each principle. Also, to test homophily, they consider hobby, favorite color, and location, while we explore key demographic features.
@@
Using LLMs in social contexts raises concerns of biases and stereotyping (Cheng, Durmus, and Jurafsky 2023; Cheng, Piccardi, and Yang 2023; Wang, Morgenstern, and Dickerson 2024). When responding to public opinion or political questions, LLMs’ answers often skew liberal (Santurkar et al. 2023; Hartmann, Schwenzow, and Witte 2023). When assigned a persona, LLMs show worse reasoning capabilities when assigned certain demographics (Gupta et al. 2024) and produce more toxic content under certain personas (Deshpande et al. 2023). However, bias in the context of social network generation remains unexplored. In this work, we investigate such biases by studying the effects of demographic variables on LLM-generated social networks.
Figure 1: Our three prompting methods to generate social networks with LLMs. See full prompts in Figures C3-C5.
@@
Deep learning approaches aim to learn graph generation directly from observed data, resulting in realistic generated networks (You et al. 2018; Simonovsky and Komodakis 2018; Guo and Zhao 2023). However, these models require a set of training graphs, and typically have fixed feature sets, so they cannot be applied in situations without access to real-world network data, and cannot be easily extended to incorporate new features, while LLMs can be. In contrast, some simple classical models can generate networks with only 1-2 parameters, which means they can generate networks without requiring training and with minimal need to define parameters. However, their simplicity often results in unrealistic network structures. For example, Erdős–Rényi models assume that each edge forms with uniform probability (Erdős and Rényi 1959), Watts–Strogatz models (Watts and Strogatz 1998) generate small-world networks but struggle to produce realistic degree distributions, and Barabási–Albert models (Barabási and Albert 1999) capture power-law degree distributions but miss community structure and clustering.
@@
The use of LLMs has been explored for various graph tasks (Li et al. 2024; Jin et al. 2024), such as graph reasoning (Wang et al. 2023; Fatemi, Halcrow, and Perozzi 2024), node classification (Zhao et al. 2023; Chen et al. 2024; Ye et al. 2024), or tasks on knowledge graphs (Pan et al. 2024). These works have shown that LLMs possess preliminary graph reasoning capabilities, but struggle on larger graphs or harder tasks, such as finding a Hamilton path (Wang et al. 2023). Yao et al. (2024) is one of the first to explore LLMs for graph generation, but they explore it in the context of molecule generation, not social networks.
@@ 3 Generating Social Networks
Our process for generating social networks involves two steps: first, constructing a set of personas, and second, having the LLM generate networks over those personas. Here, we summarize the two steps, with details in Appendix C.
@@ 3.1 Persona construction
For each persona, we include their gender, age, race/ethnicity, religion, and political affiliation, which are salient dimensions of homophily in real social networks (McPherson, Smith-Lovin, and Cook 2001; Halberstam and Knight 2016). We sample these characteristics based on the distribution of the US population. Using US Census data (US Census Bureau 2023), we acquire the joint distribution for gender, age, and race/ethnicity. Then, we sample the persona’s religion, conditioned on their race/ethnicity (Statista 2016; PRRI Staff 2021), and political affiliation, conditioned on their gender and race/ethnicity (Pew Research Center 2024; Sanchez and Foxworth 2022). In Section 5.3, we also experiment with including interests for each persona, instead of only demographic variables.
@@ 3.2 Network generation
In our first method, which we call “Global”, we provide the LLM with the entire list of personas, and prompt it to construct the network between them, in the form of edge pairs (referring to each persona’s ID).
@@
In our second method, which we call “Local”, we have the LLM take on the identity of one persona at a time, e.g., by saying, “You are a Woman, Asian, age 54, Hindu, Independent.” We provide the LLM with the list of all other personas (in the same format) and prompt it to pick friends for the persona it is currently assigned. To construct the entire network, we iterate through all personas in a random order, and we keep an edge between personas A and B if the LLM selects B when acting as A or vice versa (so we do not require both to select each other). This method is inspired by techniques in machine learning that similarly model the graph generation process by iterating through nodes and selecting edges for each node at a time (You et al. 2018).
@@
In our third method, we also assign the LLM one persona at a time, but in addition to providing the list of all other personas, we also provide information about the constructed graph so far. We experiment with providing each persona’s full list of current friends versus only their degree (i.e., total number of friends). These variations are similar to the preferential attachment experiments in Papachristou and Yuan (2024), where they also experimented with providing neighborhood information versus only degree, although their experiments only considered the network and no demographic features. In contrast, our experiments—and in particular, the comparison of the Local to Sequential methods—reveal how providing demographic and network information compare to only providing demographic information.
@@ 4 Comparison to Real Networks
To evaluate the realism of our generated networks, we gather a set of real social networks from the CASOS (CASOS 2024) and KONECT (KONECT 2024) repositories. We kept networks that described friendships between individuals, which filtered out other types of networks, such as work-related interactions or visiting ties between families. We included eight real networks, which capture friendships within diverse communities, such as among physicians, students, and prisoners (see Appendix C.3 for details).
We extract graph-level and node-level metrics from the real networks and our generated networks, and compare their distributions. For consistency, we treat all networks as undirected. Since the number of nodes varies across networks, we focus on network metrics that are comparable across graphs of different sizes, and scale those that are dependent on network size based on how they are expected to scale in an Erdös-Rényi random graph (Erdős and Rényi 1959). Below, we define and motivate the network metrics that we evaluate on.
@@
A basic property of a network is its density of edges, and social networks tend to be sparse, meaning lower density (Wong, Pattison, and Robins 2006). Density computed as the number of observed edges divided by the total number of possible edges in the network, which comes out to $\frac{2E}{N(N-1)}$, where $N$ is the number of nodes and $E$ is the number of edges in the network.
@@
Social networks are known to exhibit clustering, where one’s friends are likely to be friends with each other (Alizadeh 2017). For a node $i$, its clustering coefficient is $\frac{2E_{i}}{N_{i}(N_{i}-1)}$, where $N_{i}$ is its number of neighbors and $E_{i}$ is the number of edges between its neighbors. The average clustering coefficient computes the average over nodes.
Largest connected component (LCC).
Social networks are known to be well-connected (Ugander et al. 2011), with the vast majority (over 99%) of the nodes in the largest connected component (LCC), i.e., the largest subgraph where all nodes within the subgraph are reachable by each other. Thus, as a metric, we compute the proportion of all nodes in the network that are in the LCC, $\frac{N_{LCC}}{N}$.
Average shortest path.
Social networks are not only well-connected, meaning nodes can reach each other, but also they can reach each other in relatively short paths (Alizadeh 2017). So, we measure the average shortest path over all pairs of nodes in the LCC, divided by $\log N_{LCC}$, since shortest paths scale with $\log N$ in Erdös-Rényi graphs (Watts and Strogatz 1998). We compute shortest paths within the LCC instead of the entire network, since the distance between two disconnected nodes is infinite.
@@
Social networks exhibit strong community structure, with more edges within communities and fewer edges across communities (Newman 2004). To measure community structure, first we use the Louvain algorithm (Blondel et al. 2008) to partition the network into communities, then we assess the quality of the partition with modularity (Eq. 5). Higher levels of community structure correspond to higher modularity.
@@
Social networks are said to follow a power law degree distribution, where $P(k)\sim k^{-\gamma}$, for degree $k$ and constant $\gamma$ (Barabási and Albert 1999). This results in long-tailed degree distributions with a few people having far more friends than most others. To measure degree distribution, we compute the degree of each node in the network, and to make degree comparable across graphs, we divide degree by $N$. To summarize degree distribution from a set of networks, we pool all of the (normalized) degrees of nodes in the networks in the set, and compute the distribution over the pooled degrees in bins of 0.05, from 0 to 1.
@@
Finally, social networks are known to exhibit homophily, where “birds of a feather flock together”, i.e., people with similar traits are likelier to be friends (McPherson, Smith-Lovin, and Cook 2001). To measure homophily, we use a common metric adopted in prior work (McPherson, Smith-Lovin, and Cook 2001; Easley and Kleinberg 2010; Smith, McPherson, and Smith-Lovin 2014): the ratio of observed-to-expected cross-group edges. Specifically, first we compute the observed proportion of edges that are cross-group (e.g., different gender), then we compute the expected proportion of edges that are cross-group (based on the number of nodes that belong to each group), then we take the ratio of these two proportions. We define this as
$H$ $=\frac{C_{\textrm{obs}}}{C_{\textrm{exp}}}=\frac{\frac{\sum_{ij}A_{ij}\cdot\mathds{1}[g_{i}\neq g_{j}]}{E}}{\frac{\sum_{g}\sum_{g^{\prime}\neq g}N_{g}N_{g^{\prime}}}{N(N-1)}},$ (1)
where $A_{ij}$, as the adjacency matrix, is 1 if nodes $i$ and $j$ are connected and 0 otherwise; $g_{i}$ indicates node $i$’s group; and $N_{g}$ is the number of nodes in group $g$ [^1] For age, we use the average age difference in observed edges divided by the expected age difference (i.e., average difference over all possible pairs of nodes). If the ratio is below 1, this indicates homophily, since there are fewer cross-group edges than expected, while ratios above 1 indicate heterophily (e.g., this appears in heterosexual dating networks).